OpenStudy (anonymous):

log_2 ⁡(x-1) ≤ log_4⁡ (3-x) ...

OpenStudy (anonymous):

there may be a snap way to do this, but i think it requires the following work: first rewrite using the base of 2 using the change of base formula to get \[\log_2(x-1)<\frac{\log_2(3-x)}{\log_2(4)}=\frac{\log_2(3-x)}{2}\] \[2\log(x-1)<\log(3-x)\] \[log((x-1)^2)<\log(3-x)\] and since log base 2 is strictly increasing this means \[(x-1)^2<3-x\]

OpenStudy (anonymous):

then you still have to solve the quadratic inequality \[x^2-2x+1\leq 3-x\] \[x^2-x-2\leq 0\] \[(x-2)(x+1)\leq 0\] \[-1\leq x\leq 2\] and then you STILL have to check that the solutions are in the domain of your original inequality

OpenStudy (anonymous):

domain of \[\log_2(x-1)\] is \[x>1\] and domain of \[\log_4(3-x)\] is \[x<3\] so your actual solution is the interval \[(1,3)\] or \[1<x<3\]

OpenStudy (anonymous):

i got it... i just did not know how ti speculate this one.. no idea how to get to next sequence \[\log_{2} x-1 = 1/2 \log_{2} 3-x\].. thank you

OpenStudy (anonymous):

oh wait i made a mistake!

OpenStudy (anonymous):

solution is \[1<x\leq2\] sorry